An AC circuit is powering an electric heater (i.e., pure resistance, PF = 1.0). Assume

the voltage is 120 V and the current draw is 6 A. Compute the apparent power and

real power
Answer #1

given voltage = 120 V

Current = 6 A

power factor =1

cos\phi=1

Apparent power S=V*I

S=120*6

S= 720 VA

Real power P= V*I*cos\phi

P= 120*6*1

P=720 w.